OpenStudy (anonymous):

Using fundamental identities how do you writw the first expression in terms of the second for any acute angle : Sec of theta, Sin of theta

OpenStudy (anonymous):

cot = cos/sin. cos = cos/sin * sin = cot * sin

OpenStudy (anonymous):

i found this idk if thats what your looking for or not

OpenStudy (anonymous):

sec = 1/cos To get it in terms of sin, multiply the right side by sin/sin, giving sec = sin/(sin)(cos) Recognize that sin/cos = tan. Therefore, sec = tan/sin

Directrix (directrix):

Refresh your reciprocals. Chart attached.